Hi. Can you please explain to me how to work out decimal numbers. Im trying to learn for a test next tuesday.
I have a sample question .03 x .50. Im unsure how to work out the answer.
Thanks
Brittany

Answer by mananth(16946)

I have a sample question .03 x .50. Im unsure how to work out the answer.
When there are decimals students get confused. This is very simple.
when there are decimals just carry out the multiplication as if there were no decimals.
For example
1.5 * 1.5
you know 15 * 15 = 225
count the number of digits after the decimal point in both the numbers.
1+1 =2
so just put a decimal point after two places counting from the units place.
1.5*1.5 = 2.25
..
2.24 *0.6
without decimal 1344
count the number of digits after the decimal point in both the numbers.
3
count from units place 3 digits & put the decimal point.
...
1.25 * 0.004
same method
without decimal 125 * 4 = 500
count the number of digits after the decimal point in both the numbers.
total decimals = 2+3=5
0.00500
...
1.62 * 0.00002
162*2=324
total decimals = 2+5=7
0.0000324
.
m.ananth@hotmail.ca